OpenStudy (anonymous):

Given the following perfect square trinomial, fill in the missing term. x^2 - 16x + ____

OpenStudy (anonymous):

64

OpenStudy (anonymous):

64

OpenStudy (anonymous):

divide -16 by 2 = -8 and then square it (-8)^2 = 64

OpenStudy (hoblos):

in general it is a^2 - 2ab + b^2 a=x -2ab=-16x so b=8 b^2 = 64